ISOTOPOMER DEPENDENCE AND BORN-OPPENHEIMER BREAKDOWN IN MID AND FAR IN-FRARED SPECTRA OF CdH

\underline{ROBERT J. LE ROY}, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; THOMAS D. VARBERG, Department of Chemistry, Macalester College, 1600 Grand Ave., St. Paul, MN 55105, USA.

New measurements of far infrared transitions within the v =1 levels of the X^{2}Σ^{+} state of 6 different isotopomers of CdH and CdD are combined with previous mid-infrared^{a}^{,}^{b} and far-infrared^{c} data in a combined-isotopomer analysis which yields improved rotational constants, spin-rotation constants and Born-Oppenheimer breakdown (BOB) parameters for this system. Recent work by one of us^{c} had reported the measurement and analysis of far infrared transitions within the v = 0 level of 12 isotopomers CdH and CdD, but the combined-isotopomer analysis reported there found that the CdH and CdD data appeared to be incompatible unless separate Cd-atom BOB parameters were introduced for the inertial rotation constants of CdH and CdD, and BOB parameters introduced for the spin-rotation constants. The present work shows that taking account of the v–dependence of the mechanical rotationl parameters for v = 0 isotopomers implied by the reduced-mass-scaling of vibrational quantum numbers allows the far infrared data to be accounted for without invoking these two assumptions.

a R.-D. Urban, U. Magg, H. Birk and H. Jones, J. Chem. Phys. 92, 14-21 (1990).

b H. Birk, R.-D. Urban, P. Polomsky and H. Jones, J. Chem. Phys. 94, 5435-5442 (1991).

c T.D. Varberg and J.C. Roberts, J. Mol. Spectrosc. 223, 1-8 (2004)